### v1

# A note on division rings satisfying generalized rational identities with anti-automorphisms

Abstract.
Let $D$ be a division ring with infinite center $F$; $\sigma$ be an anti-automorphism of $D$ and $m$ be a positive integer such that $\sigma^{m}\neq\mathrm{Id}$. In this paper, we show that if $D$ satisfies a $\sigma^{m}$-GRI, then $D$ is centrally finite.
Key words and phrases:
anti-automorphism, division ring, rational identity.
2020 Mathematics Subject Classification. 16K20, 16R50
1Faculty of Mathematics and Computer Science, University of Science,
Ho Chi Minh City, Vietnam
2Vietnam National University, Ho Chi Minh City, Vietnam
3Binh Duong University, Binh Duong Province, Vietnam
minhthuvohoang@gmail.com, vmtrang@bdu.edu.vn
1. Introduction and the main result
Let $D$ be a division ring with center $F$. We say that $D$ is centrally finite if $D$ is regarded as a finite-dimensional vector space over $F$. This topic relates to conditions under which a division ring is centrally finite. We focus on division rings satisfying certain identities.
A classic result states that if $D$ satisfies the polynomial identity, then $D$ is centrally finite (see [9]). To present this result and the extended results we aim to prove, we recall some concepts. Let $X=\{x_{1},x_{2},\cdots,x_{n}\}$ be a set of $n$ non-commutative indeterminates and $K$ be a field. We denote $K\langle X\rangle$ as the polynomial ring in indeterminates $X$ over $K$, and $R$ as a ring whose center contains $K$. A polynomial $f(X)\in K\langle X\rangle\backslash\{0\}$ is called a polynomial identity of $R$ if for every $n$-tuple $(a_{1},a_{2},\dots,a_{n})\in R^{n}$, we have $f(a_{1},a_{2},\dots,a_{n})=0$. In this case, we say $R$ satisfies the polynomial identity $f$. If $R$ satisfies some polynomial identity, then we say $R$ satisfies a polynomial identity. Since we only work with division rings, we directly assume $R=D$ is a division ring and $K$ coincides with the center $F$ of $D$.
The most classical result concerning division rings satisfying polynomial identities relates to its finite-dimensionality. Specifically, it is proven in [9] that if $D$ is a division ring with center $F$ and satisfies some polynomial identity over $F$, then $D$ is centrally finite. This result has been subsequently extended in many different directions (e.g., see [1, 2, 3, 5, 12, 10, 11]). More precisely, we denote $D\langle X\rangle=D*_{F}F\langle X\rangle$ as the free product of $D$ with $F\langle X\rangle$ over $F$. Each element in $D\langle X\rangle\backslash\{0\}$ is called a generalized polynomial over $D$. Furthermore, since the ring of generalized polynomials $D\langle X\rangle$ is fir, the maximal right ring of quotients of $D\langle X\rangle$ is a division ring (e.g., see [6]). This division ring is denoted as $D(X)$. Each nonzero element in $D(X)$ is called a generalized rational polynomial over $D$. Suppose $f(X)$ is a generalized rational polynomial (respectively, a generalized polynomial) over $D$. If for every $n$-tuple $(a_{1},a_{2},\dots,a_{n})\in D^{n}$ such that when $f(a_{1},a_{2},\dots,a_{n})$ is defined, one has $f(a_{1},a_{2},\dots,a_{n})=0$, then we say $D$ satisfies the generalized rational identity (respectively, the generalized identity) $f$, or $f$ is a generalized rational identity (respectively, a generalized identity) of $D$.
In 1966, Amitsur proved that if the center $F$ is infinite and the division ring $D$ satisfies a generalized rational identity, then $D$ is centrally finite (see [2]). There are several extensions of this result. For example, in 1996, Chiba proved that if the center $F$ is infinite and the multiplicative group $D^{*}=D\backslash\{0\}$ contains a non-central proper subnormal subgroup satisfying a generalized rational identity, then $D$ is centrally finite (see [4]). Another direction is to consider identities corresponding to an anti-automorphism. Recall that a bijection $\sigma:D\to D,x\mapsto x^{\sigma},$ is called an anti-automorphism of $D$ if $(a+b)^{\sigma}=a^{\sigma}+b^{\sigma}$ and $(ab)^{\sigma}=b^{\sigma}a^{\sigma}$ for all $a,b\in D$. In the case where the order of $\sigma$ is $2$, $\sigma$ is called an involution of $D$. Let $m$ be a positive integer such that $\sigma^{i}\neq\mathrm{Id}$ for all $1\leq i\leq m$. For each $1\leq i\leq m$, denote $X^{\sigma^{i}}=\{x_{1}^{\sigma^{i}},x_{2}^{\sigma^{i}},\dots\}$ as the set of indeterminates indexed corresponding to each $\sigma^{i}$ and $X_{m}=\cup_{i=0}^{m}X^{\sigma^{i}}$. We denote $D(X_{m})$ as the division ring as mentioned above. Let $f(X_{m})$ be a nonzero element in $D(X_{m})$. If for every $n$-tuple $(r_{1},r_{2},\cdots,r_{n})\in D^{n}$ such that when
$f(r_{1},r_{2},\dots,r_{n},r_{1}^{\sigma},r_{2}^{\sigma},\dots,r_{n}^{\sigma},\dots,r_{1}^{\sigma^{m}},r_{2}^{\sigma^{m}},\dots,r_{n}^{\sigma^{m}})$
is defined, we have
$f(r_{1},r_{2},\dots,r_{n},r_{1}^{\sigma},r_{2}^{\sigma},\dots,r_{n}^{\sigma},\dots,r_{1}^{\sigma^{m}},r_{2}^{\sigma^{m}},\dots,r_{n}^{\sigma^{m}})=0,$
then we say $f(X_{m})$ is a generalized rational identity of $D$ with respect to $\sigma^{m}$. In this case, we simply say that $D$ satisfies a $\sigma^{m}$- GRI $f$ or $f$ is a $\sigma^{m}$-GRI of $D$. The ”best” result is related to when $\sigma$ is an involution. In this case, $m=1$ and $\sigma$ will be denoted as $\star$. In more detail, it is proved that if the center $F$ is infinite and $D$ satisfies a $\star$-GRI, then $D$ is centrally finite (see [7]). The case of generalized polynomial identities with respect $\sigma$ receives even more attention (e.g., see the book [3] for a survey on this topic). The objective of this paper is to prove the following result.
Theorem 1.1.
Let $D$ be a division ring with infinite center $F$. If $D$ satisfies a $\sigma^{m}$-GRI then $D$ is centrally finite.
This theorem can be viewed as an anti-automorphism counterpart of Amitsur’s Theorem and a natural extension of [7].
2. Preliminaries
In short and simple terms, we prove the main result of this paper by induction on $n$, the number of indeterminates in the identity. In this section, we prove the case for $n=1$. To do it, we need to review some basic concepts and establish notation. Many results in this section may have proofs similar to those found in some reference documents, but none of the results are directly quoted from them. Therefore, for the convenience of the readers, we will provide proofs for the results if necessary. The main result we obtained in this section is Lemma  2.4, which states that if a division ring satisfies a homogeneous linear identity in one variable, then this division ring is centrally finite.
We emphasize some notations that will be used throughout the paper as follows. Let $D$ be a division ring with center $F$, and $\sigma:D\to D,x\mapsto x^{\sigma},$ be an anti-automorphism over $F$, that is, $(a+b)^{\sigma}=a^{\sigma}+b^{\sigma},(ab)^{\sigma}=b^{\sigma}a^{\sigma}$ and $\alpha^{\sigma}=\alpha$ for every $a,b\in D$ and $\alpha\in F$. Consider $X=\{x_{1},x_{2},\dots\}$ as a countable set of non-commutative indeterminates. Suppose $m$ is an integer such that $\sigma^{i}\neq Id$ for all $1\leq i\leq m$. For each $1\leq i\leq m$, let $X^{\sigma^{i}}=\{x_{1}^{\sigma^{i}},x_{2}^{\sigma^{i}},\dots\}$ and $X_{m}=\bigcup_{i=0}^{m}X^{\sigma^{i}}$. The notation $F\langle X_{m}\rangle$ denotes the free algebra generated by $X_{m}$ over $F$, and $D\langle X_{m}\rangle=D*_{F}F\langle X_{m}\rangle$ is the free product of $D$ and $F\langle X_{m}\rangle$ over $F$. Each element of $D\langle X_{m}\rangle$ has the form
$f=f(x_{1},x_{2},\dots,x_{n},x_{1}^{\sigma},x_{2}^{\sigma},\dots,x_{n}^{\sigma},\dots,x_{1}^{\sigma^{m}},x_{2}^{\sigma^{m}},\dots,x_{n}^{\sigma^{m}})=\sum_{i=1}^{l}P_{i},$
where $P_{i}=\delta_{i}a_{i_{0}}x_{i_{1}}^{\sigma^{k_{1}}}a_{i_{1}}x_{i_{2}}^{\sigma^{k_{2}}}a_{i_{2}}\cdots x_{i_{q}}^{\sigma^{k_{q}}}a_{i_{q}}$ are monomials, $\delta_{i}$ is in $F$, $\{a_{i_{0}},a_{i_{1}},\dots,a_{i_{q}}\}$ is a subset of $D$, and $\{x_{i_{1}}^{\sigma^{k_{1}}},x_{i_{2}}^{\sigma^{k_{2}}},\dots,x_{i_{q}}^{\sigma^{k_{q}}}\}$ is a subset $X_{m}$.
For each monomial $P\in D\langle X_{m}\rangle$, we define the $\sigma^{m}$-degree of $P$ with respect to the indeterminate $x_{j}$, denoted $\sigma^{m}$-$\deg_{x_{j}}(P)$, as the number of occurrences of $x_{j}^{\sigma^{i}}$ in $P$ for some $0\leq i\leq m$. Then, the degree of $P$, denoted $\deg(P)$, is defined as
$\deg(P)=\sum_{x_{j}\in X}\sigma^{m}\text{-}\deg_{x_{j}}(P).$
We also define the $\sigma^{m}$-height of $P$ with respect to the indeterminate $x_{j}$, denoted $\sigma^{m}\text{-}ht_{x_{j}}(P)$, as the maximum value between $(\sigma^{m}\text{-}\deg_{x_{j}}(P)-1)$ and 00. Similarly, the height of $P$, denoted $ht(P)$, is defined as
$ht(P)=\sum_{x_{j}\in X}\sigma^{m}\text{-}ht_{x_{j}}(P).$
In other words, the height of $P$ is equal to the degree of $P$ minus the number of distinct indeterminates appearing in $P$. Let $\AA$ be a $F$-basis of $D$ containing $1$. Then $\AA$, combined with the standard $F$-basis of $F\langle X_{m}\rangle$, induces a monomial basis $M(\AA)$ for $D\langle X_{m}\rangle$. Suppose $f=\sum_{i=1}^{\ell}P_{i}\in D\langle X_{m}\rangle$ where $P_{1},P_{2},\dots,P_{\ell}$ are monomials in $M(\AA)$. We also define
  1. (1)
    $\sigma^{m}\text{-}\deg_{x_{j}}(f)=\max\{\sigma^{m}\text{-}\deg_{x_{j}}(P_{i})\mid 1\leq i\leq l\}$;
  2. (2)
    $\deg(f)=\max\{\deg(P_{i})\mid 1\leq i\leq l\}$;
  3. (3)
    $\sigma^{m}\text{-}ht_{x_{j}}(f)=\max\{\sigma^{m}\text{-}ht_{x_{j}}(P_{i})\mid 1\leq i\leq l\}$;
  4. (4)
    $ht(f)=\max\{ht(P_{i})\mid 1\leq i\leq l\}$;
Assume that $x_{1},x_{2},\dots,x_{n}$ are indeterminates appearing in the polynomial $f$. Suppose indeterminate $x_{j}$ appears in all monomials of $f$, then we say $f$ is homogeneous with respect to the indeterminate $x_{j}$. If $f$ is homogeneous with respect to every indeterminate $x_{j}$ where $1\leq j\leq n$, then we say $f$ is homogeneous.
Suppose $f=f(x_{1},x_{2},\dots,x_{n},x_{1}^{\sigma},x_{2}^{\sigma},\dots,x_{n}^{\sigma},\dots,x_{1}^{\sigma^{m}},x_{2}^{\sigma^{m}},\dots,x_{n}^{\sigma^{m}})$ is a $\sigma^{m}$-GPI of $D$ with $\deg(f)=p$. Let $q$ be the number of indeterminates in $f$ that are not homogeneous. Suppose $q>0$. Without loss of generality, we assume $f$ is not homogeneous with respect to indeterminates $x_{1},x_{2},\cdots,x_{q}$. Set
$f_{1}$ $=f(0,x_{2},\dots,x_{n},0,x_{2}^{\sigma},\dots,x_{n}^{\sigma},\dots,0,x_{2}^{\sigma^{m}},\dots,x_{n}^{\sigma^{m}});$
$f_{2}$ $=f-f_{1}.$
Since $f$ is a $\sigma^{m}$-GPI of $D$, one has $f_{1}$ and $f_{2}$ are also $\sigma^{m}$-GPIs of $D$. It is clear that $\deg(f_{1})$ and $\deg(f_{2})$ are less than or equal to $p$, and $f_{2}$ is homogeneous with respect to indeterminate $x_{1}$. Similarly, for $2\leq j\leq q$, set
$f_{2j-1}$ $=f_{2j-2}(x_{1},x_{2},\dots,x_{j-1},0,$
$\quad x_{j+1},\cdots,x_{n},\cdots,x_{1}^{\sigma^{m}},x_{2}^{\sigma^{m}},\dots,x_{j-1}^{\sigma^{m}},0,x_{j+1}^{\sigma^{m}},\cdots,x_{n}^{\sigma^{m}}),$
and $f_{2j}=f_{2j-2}-f_{2j-1}$. Then, the polynomial $f_{2q}$ is homogeneous with respect to indeterminates $x_{1},x_{2},\cdots,x_{n}$, and is also a $\sigma^{m}$-GPI of $D$. From the above reasoning, we obtain the following proposition.
Proposition 2.1.
Let $D$ be a division ring and $\sigma$ be an anti-automorphism of $D$. If $D$ satisfies a $\sigma^{m}$-GPI, then $D$ satisfies a $\sigma^{m}$-GPI that is homogeneous.
We continue to consider $f=\sum_{i=1}^{\ell}P_{i}\in D\langle X_{m}\rangle$ where $P_{1},P_{2},\dots,P_{\ell}$ are monomials. Then, $f$ is called $\sigma^{m}$-multilinear (or shortly, $\sigma^{m}$-linear) of degree $n$ on indeterminates $x_{1},x_{2},\cdots,x_{n}$ if each monomial $P_{i}$ in $f$ satisfies the following properties:
  1. (1)
    $\deg(P_{i})=n$;
  2. (2)
    $\sigma^{m}\text{-}\deg_{x_{j}}(P_{i})=1$ for all $x_{j}\in\{x_{1},x_{2},\cdots,x_{n}\}$.
In other words, the polynomial $f$ is $\sigma^{m}$-linear if $f$ is a polynomial homogeneous of height 00.
Suppose $f$ is a homogeneous $\sigma^{m}$-GPI of $D$ with $\{x_{1},x_{2},\dots,x_{n}\}$ being the set of all indeterminates appearing in $f$. Let $p=\deg(f)$ and $p_{j}=\sigma^{m}\text{-}\deg_{x_{j}}(f)$ $(1\leq j\leq k)$. We write $f=f(x_{j})$ to emphasize considering $f$ as a polynomial in the indeterminate $x_{j}$. If $ht(f)=t>0$, we choose $x_{j_{q}}\in X$ such that $x_{j_{q}}$ does not appear in $f$ and set
$\Delta_{j}f=f(x_{j}+x_{j_{q}})-f(x_{j})-f(x_{j_{q}}).$
Then:
  1. (1)
    $\deg(\Delta_{j}f)=\deg(f)$;
  2. (2)
    $\sigma^{m}\text{-}\deg_{x_{j}}(\Delta_{j}f)=\sigma^{m}\text{-}\deg_{x_{j_{q}}}(\Delta_{j}f)=p_{j}-1$;
  3. (3)
    $\sigma^{m}\text{-}\deg_{x_{j_{k}}}(\Delta_{j}f)\leq\sigma^{m}\text{-}\deg_{x_{j_{k}}}(f)$ for $x_{j_{k}}\neq x_{j},x_{j_{q}}$.
Therefore, $\Delta_{j}f$ is also a homogeneous $\sigma^{m}$-GPI of $D$ satisfying $ht(\Delta_{j}f)\leq t-1$. Define
$\Delta_{j}^{(0)}f$ $=f,$
$\Delta_{j}^{(1)}f$ $=\Delta_{j}f,$
$\Delta_{j}^{(v)}f$ $=\Delta_{j}\Delta_{j}^{(v-1)}f\quad(v\in\mathbb{N},v>1).$
Then, if $\sigma^{m}\text{-}\deg_{x_{j}}(f)=p_{j}$, we have $\sigma^{m}\text{-}\deg_{x_{j}}(\Delta_{j}^{(p_{j}-1)}f)=1$. From here, we have the following lemma.
Lemma 2.2.
Given $f=\sum_{i=1}^{\ell}P_{i}\in D\langle X_{m}\rangle$ is a homogeneous polynomial in indeterminates $x_{1},x_{2},\dots,x_{n}$ and has height $t$. Let $P$ be a monomial in $f$ such that $ht(P)=t$. Then, $\Delta_{1}^{(p_{1}-1)}\dots\Delta_{n}^{(p_{n}-1)}f$ is a $\sigma^{m}$-linear polynomial in $\sum_{j=1}^{n}p_{j}$ indeterminates, where $p_{j}=\sigma^{m}\text{-}\deg_{x_{j}}(P)$ $(1\leq j\leq n)$.
Proof.  If $f$ has height $t=0$, then $p_{j}=1$ and $\Delta^{(0)}_{j}f=f$ is $\sigma^{m}$-linear for every $1\leq j\leq n$, thus the lemma is proved. Consider the case $t>0$ and $P$ is a monomial in $f$ such that $\sigma^{m}\text{-ht}(P)=t$. With $p_{j}=\sigma^{m}\text{-}\deg_{x_{j}}(P)$, we have
$\Delta_{1}^{(p_{1}-1)}\dots\Delta_{n}^{(p_{n}-1)}f=\sum\Delta_{1}^{(p_{1}-1)}\dots\Delta_{n}^{(p_{n}-1)}P_{i}.$
Choose $j_{0}$ such that $\sigma^{m}\text{-}\deg_{x_{j_{0}}}(P)=p_{j_{0}}>1$ and $\deg_{x_{i}}(P)=p_{i}\leq 1$ in which $1\leq j_{0}<i\leq n$. Then
$\Delta_{1}^{(p_{1}-1)}\dots\Delta_{n}^{(p_{n}-1)}P=\Delta_{1}^{(p_{1}-1)}\dots\Delta_{j_{0}}^{(p_{j_{0}}-1)}P.$
Let $x_{n+1}\in X$, set
$\Delta_{j_{0}}P=P(x_{j_{0}}+x_{n+1})-P(x_{j_{0}})-P(x_{n+1})={\sum_{u=1}^{p_{j_{0}}-1}h_{u}}$
in which $h_{u}$ is a polynomial such that $\sigma^{m}\text{-}\deg_{x_{n+1}}(h_{u})=u$ and $\sigma^{m}\text{-}\deg_{x_{j_{0}}}(h_{u})=p_{j_{0}}-u$. Each polynomial $h_{u}$ is a sum of $C_{p_{j_{0}}}^{u}$ monomials $P_{u_{k}}$, where each monomial $P_{u_{k}}$ maps to $P$ under the projection $x_{n+1}\to x_{j_{0}}$ and $\sigma^{m}\text{-}\deg_{x_{j_{0}}}(P_{u_{k}})=p_{j_{0}}-u$. Hence,
$\Delta_{j_{0}}^{(p_{j_{0}}-1)}P=\Delta_{j_{0}}^{(p_{j_{0}}-2)}\Delta_{j_{0}}^{(1)}f=\Delta_{j_{0}}^{(p_{j_{0}}-2)}\sum_{u=1}^{p_{j_{0}}-1}h_{u}=\sum_{u=1}^{p_{j_{0}}-1}\Delta_{j_{0}}^{(p_{j_{0}}-2)}h_{u}.$
For $u>1$, since $\sigma^{m}$-$\deg_{x_{j_{0}}}(h_{u})=p_{j_{0}}-u<p_{j_{0}}-1$, we have $\Delta_{{{j_{0}}}}^{(p_{j_{0}}-2)}(h_{u})=0$. Therefore,
$\Delta_{{j_{0}}}^{p_{j_{0}}-1}P=\Delta^{(p_{j_{0}}-2)}_{{j_{0}}}h_{1}=\Delta^{(p_{j_{0}}-2)}_{{j_{0}}}\left(\sum^{C^{1}_{p_{j_{0}}}}_{k=1}P_{1_{k}}\right)=\sum^{C^{1}_{p_{j_{0}}}}_{k=1}\Delta^{(p_{j_{0}}-2)}_{j_{0}}P_{1_{k}}.$
For each $1\leq k\leq C^{1}_{p_{j_{0}}},$ we have
$\Delta^{(p_{j_{0}}-2)}_{{j_{0}}}P_{1_{k}}=\Delta^{(p_{j_{0}}-3)}_{j_{0}}\left(\Delta^{(1)}_{j_{0}}P_{1_{k}}\right)=\Delta^{(p_{j_{0}}-3)}_{j_{0}}\left(\sum^{p_{j_{0}}-2}_{u_{k}=1}h^{{}^{\prime}}_{u_{k}}\right)=\sum^{p_{j_{0}}-2}_{u_{k}=1}\Delta^{(p_{j_{0}}-3)}_{j_{0}}h^{{}^{\prime}}_{u_{k}}.$
On the other hand, each polynomial $h^{{}^{\prime}}_{u_{k}}$ has $\sigma^{m}$-$\deg_{x_{n+2}}(h^{{}^{\prime}}_{u_{k}})=u_{k}$ and $\sigma^{m}$-$\deg_{x_{j_{0}}}(h^{{}^{\prime}}_{u_{k}})=p_{j_{0}}-1-u_{k}$. Note that $h^{{}^{\prime}}_{u_{k}}$ has $C^{u_{k}}_{p_{j_{0}}-1}$ monomials, each of which maps to $P$ under the projection $x_{j_{k}}\mapsto x_{j_{0}}$ (where $j_{k}=n+1,n+2$). Similarly, for each $u_{k}>1$, we also have $\Delta_{j}^{(p_{j}-3)}h^{{}^{\prime}}_{u_{k}}=0$. Therefore, $\Delta_{j_{0}}^{(p_{j_{0}}-2)}P_{1_{k}}=\Delta_{j_{0}}^{(p_{j}-3)}h^{{}^{\prime}}_{1}$. Thus, $\Delta_{j_{0}}^{(p_{j_{0}}-1)}P=\sum_{u=1}^{C^{p_{j_{0}}}_{1}}\sum^{C^{p_{j_{0}}-1}}_{v=1}\Delta_{{j_{0}}}^{p_{j_{0}}-3}h^{"}_{uv}$, where $h"_{uv}$ are monomials. Continuing these steps, we see that $\Delta_{j}^{p_{j}-1}P$ has
$C^{1}_{p_{j}}C^{1}_{p_{j}-1}\cdots C^{1}_{2}=p_{j}(p_{j}-1)\cdots 2=p_{j}!$
monomials, each of which maps to $P$ under the projection $x_{e}\mapsto x_{j_{0}}$ (and here $e=n+1,n+2\cdots,n+(p_{j_{0}}-1$). Furthermore, we have
$\sigma^{m}\text{-}\deg_{x_{j_{0}}}(\Delta_{j_{0}}^{(p_{j_{0}}-1)}P)=\sigma^{m}\text{-}\deg_{x_{e}}(\Delta_{x_{e}}^{(p_{j_{0}}-1)}P)=1$
for $e\in\{n+1,n+2,\cdots,n+(p_{j_{0}}-1)\}$. After decomposing $\Delta_{j_{0}}^{(p_{j_{0}}-1)}P$ into monomials as above, we continue this technique for $\Delta_{j_{0}-1}^{(p_{j_{0}-1}-1)}\Delta_{j_{0}}^{(p_{j_{0}}-1)}P$, and further expand $\Delta_{j_{0}-2}^{(p_{j_{0}-2}-1)}\Delta_{j_{0}-1}^{(p_{j_{0}-1}-1)}\Delta_{j_{0}}^{(p_{j_{0}}-1)}P$. Eventually, we obtain $\Delta^{(p_{1}-1)}_{{1}}\cdots\Delta^{(p_{n}-1)}_{n}P$ as a linear polynomial in $\sum^{n}_{j=1}p_{j}$ indeterminates. Moreover, these monomials all map to $P$ under the projection satisfying the property: for every $1\leq j\leq n$, the projection of $x_{e}$ to $x_{j}$ is such that $j+\sum_{u=j+1}^{n}p_{u}\leq e\leq j+\sum_{u=j}^{n}p_{u}$.
Let $P_{i_{0}}\neq P$ be a monomial of $f$. If there exists $1\leq j_{k}\leq n$ such that $\sigma$-$\deg_{x_{{j_{k}}}}(P_{i_{0}})<p_{j_{k}}=\sigma$-$\deg_{x_{j_{k}}}(P)$, then $\Delta_{{j_{k}}}^{(p_{j_{k}}-1)}P_{i_{0}}=0$. So
$\Delta^{(p_{1}-1)}_{{1}}\cdots\Delta^{(p_{j_{k}-1}-1)}_{j_{k}-1}\Delta^{(p_{j_{k}}-1)}_{j_{k}+1}\Delta^{(p_{j_{k}+1}-1)}_{j_{k}+1}\cdots\Delta^{(p_{n}-1)}_{n}P_{i_{0}}\\ =\Delta^{(p_{1}-1)}_{{1}}\cdots\Delta^{(p_{j_{k}-1}-1)}_{j_{k}-1}\Delta^{(p_{j_{k}+1}-1)}_{j_{k}+1}\cdots\Delta^{(p_{l}-1)}_{x_{k_{l}}}\left(\Delta^{(p_{j_{k}}-1)}_{x_{k_{j_{k}}}}P_{i_{0}}\right)=0.$
If $\Delta^{(p_{1}-1)}_{1}\cdots\Delta^{(p_{n}-1)}_{{n}}P_{i_{0}}\neq 0$ then $p_{j}^{{}^{\prime}}=\sigma^{m}$-$\deg_{x_{j}}(P_{i_{0}})\geq p_{j},$ with $1\leq j\leq n.$ However, observe that $\sum(p_{j}-1)=t\geq\sum(p^{{}^{\prime}}_{j}-1)$, which implies that $p^{{}^{\prime}}_{j}=p_{j}$ for $1\leq j\leq n.$ Using a similar method as for $P$, we have that $\Delta^{(p_{1}-1)}_{1}\cdots\Delta^{(p_{n}-1)}_{{n}}P_{i_{0}}$ is a multilinear polynomial in $\sum p_{j}$ indeterminates just like $\Delta^{(p_{1}-1)}_{1}\cdots\Delta^{(p_{n}-1)}_{n}P$. Hence, $\Delta^{(p_{1}-1)}_{1}\cdots\Delta^{(p_{n}-1)}_{{n}}f$ is a homogeneous $\sigma^{m}$-linear polynomial.
From the lemma above, we immediately obtain the following corollary.
Corollary 2.3.
If the division ring $D$ satisfies an $\sigma^{m}$-GPI of degree $n$, then there exists an $\sigma^{m}$-GPI of $D$ which is homogeneous $\sigma^{m}$-linear of degree not exceeding $n$.
Now, we show the main result of this section by considering the case where $X=\{x\}$. In this case, if $f$ is a linear $\sigma^{m}$-form in $D\langle X_{m}\rangle$, then $f$ can be represented as $f=\sum_{i=0}^{k}g_{i}(x^{\sigma^{i}})$, where $g_{i}(x^{\sigma^{i}})=\sum_{j=1}^{n_{i}}a_{ij}x^{\sigma^{i}}b_{ij}$ and $0\leq k\leq m$.
Lemma 2.4.
Let $D$ be a division ring with an infinite center and $X=\{x\}$. Suppose $f\neq 0$ is a $\sigma^{m}$-generalized polynomial in $D\langle X_{m}\rangle$. If $D$ is not centrally finite, then there exists an element $r\in D$ such that $f(r)\neq 0$.
Proof.  By Corollary 2.3, we assume that $f$ is homogeneous $\sigma^{m}$-linear. By the hypothesis, $f$ can be represented as
$f=\sum_{i=0}^{k}g_{i}(x^{\sigma^{i}})$
where $g_{i}(x^{\sigma^{i}})=\sum_{j=1}^{n_{i}}a_{ij}x^{\sigma^{i}}b_{ij}$. We can choose the largest non-negative integer $0\leq k\leq m$ such that $g_{k}(x^{\sigma^{k}})\neq 0$. Consider the case $k=0$, then
$f=f(x)=g_{0}(x)=\sum_{j=1}^{n_{0}}a_{0j}xb_{0j}.$
Suppose $\sum_{j=1}^{n_{0}}a_{0j}rb_{0j}=0$ for all $r\in D$. With the assumption at hand, we deduce that $f(r)=\sum_{j=1}^{n_{0}}a_{0j}rb_{0j}\in DxD$ is a GPI of $D$. According to [3, Corollary 6.1.3], then $\sum_{j=1}^{n_{0}}a_{0j}xb_{0j}=0$, which contradicts the assumption. Hence, there exists an element $r\in D$ such that
$f(r)=\sum_{j=1}^{n_{0}}a_{0j}rb_{0j}\neq 0.$
Suppose the lemma holds for all $k\leq m-1$. For $k=m$, we have
$f=\sum_{i=0}^{m}g_{i}(x^{\sigma^{i}})\in D\left\langle X_{m}\right\rangle\backslash\{0\}.$
Suppose conversely, for all $r\in D$, $\sum_{i=0}^{m}g_{i}(r^{\sigma^{i}})=0.$ Consider the polynomial $g_{0}(x)=\sum^{n_{0}}_{j=1}a_{0j}xb_{0j}$ in two cases $n_{0}=0$ and $n_{0}>0$.In the case. First, we consider the case $n_{0}>0$. If the set $\{a_{01},a_{02},\cdots,a_{0n_{0}}\}$ of $n_{0}$ elements is linearly independent over $F$, then the set of $2n_{0}$ polynomials
$a_{01},a_{02},a_{03},\cdots,a_{0n_{0}},a_{mn_{n}}xa_{01},a_{mn_{m}}xa_{02},a_{mn_{m}}xa_{03},\cdots,a_{mn_{m}}xa_{0n_{0}}$
is also linearly independent over $F$. Indeed, consider
$\alpha_{1}a_{01}+\cdots+\alpha_{n}a_{0n_{0}}+\beta_{1}a_{mn_{n}}xa_{01}+\cdots+\beta_{n}a_{mn_{m}}xa_{0n_{0}}=0,$
where $\alpha_{j},\beta_{j}\in F$ and $j\in\{1;2;\cdots;n_{0}\}$. Then, we have
$\begin{cases}\alpha_{1}a_{01}+\alpha_{2}a_{02}+\cdots+\alpha_{n}a_{0n_{0}}&=0\\ \beta_{1}a_{mn_{n}}xa_{01}+\beta_{2}a_{mn_{n}}xa_{02}+\cdots+\beta_{n}a_{mn_{m}}xa_{0n_{0}}&=0.\end{cases}$
Since $\{a_{01},a_{02},\cdots,a_{0n_{0}}\}$ is linearly independent over $F$, we have the following: $\alpha_{1}=\alpha_{2}=\cdots=\ \alpha_{n}=0.$ Furthermore, as $\{a_{01},a_{02},\cdots,a_{0n_{0}}\}$ is linearly independent over $F$, the set
$\{a_{mn_{n}}\otimes_{F}a_{01};a_{mn_{m}}\otimes_{F}a_{02};\cdots;a_{mn_{m}}\otimes_{F}a_{0n_{0}}\}$
is also linearly independent over $F$. According to [3, Remark 6.1.1], the set
$\{a_{mn_{n}}xa_{01},a_{mn_{m}}xa_{02},\cdots,a_{mn_{m}}xa_{0n_{0}}\}$
is linearly independent over $F$, which implies $\beta_{1}=\beta_{2}=\cdots=\beta_{n}=0$. Thus, there exists an element $r\in D$ such that
$a_{01};a_{02};a_{03};\cdots;a_{0n_{0}};a_{mn_{n}}ra_{01};a_{mn_{m}}ra_{02};a_{mn_{m}}ra_{03};\cdots;a_{mn_{m}}ra_{0n_{0}}$
are linearly independent over $F$. Suppose these elements are linearly dependent over $F$ for all $r\in D$, meaning that there exist $\gamma_{1},\gamma_{2},\dots,\gamma_{2n}$ not all zero such that
$\gamma_{1}a_{01}+\gamma_{2}a_{02}+\dots+\gamma_{2n}a_{mn}ra_{0n_{0}}=0$
for all $r\in D$, implying that $D$ satisfies a GPI. According to [2], $D$ has a finite center, which contradicts the assumption. Let
$g=\sum_{i=0}^{m}g_{i}(\sigma^{i}(x\sigma^{-m}(a_{mn_{m}})\sigma^{-m}(r))-a_{mn_{m}}r\sum_{i=0}^{m}g_{i}(x^{\sigma^{i}}).$
Then,
$g$ $=\sum_{j=1}^{n_{0}}a_{0j}x\sigma^{-m}(a_{mn_{m}})\sigma^{-m}(r)b_{0j}-\sum_{j=1}^{n_{0}}a_{mn_{m}}ra_{0j}xb_{0j}$
$+\sum_{j=1}^{n_{1}}a_{1j}\sigma^{1-m}(r)\sigma^{1-m}(a_{mn_{m}})x^{\sigma}b_{1j}-\sum_{j=1}^{n_{1}}a_{mn_{m}}ra_{1j}xb_{1j}$
$\vdots$
$+\sum_{i=1}^{n_{m}-1}(a_{mj}ra_{mn_{m}}-a_{mn_{m}}ra_{mi})x^{\sigma^{m}}b_{mi}.$
Since
$a_{01},a_{02},a_{03},\cdots,a_{0n_{0}},a_{mn_{m}}ra_{01},a_{mn_{m}}ra_{02},a_{mn_{m}}ra_{03},\cdots,a_{mn_{m}}ra_{0n_{0}}$
are linearly independent over $F$,
$a_{01}\otimes_{F}\sigma^{-m}(a_{mn_{m}})\sigma^{-m}(r)b_{01},\dots,a_{0n_{0}}\otimes_{F}\sigma^{-m}(a_{mn_{m}})\sigma^{-m}(r)b_{0n_{0}},$
$a_{mn_{m}}ra_{01}\otimes_{F}b_{01},\dots,a_{mn_{m}}ra_{0n_{0}}\otimes_{F}b_{0n_{0}}$
are also linearly independent over $F$. According to [3, Remark 6.1.1], there exists an isomorphism $\phi:D\otimes_{F}D\to DxD$ such that $a\otimes_{F}\mapsto axb$. Since
$\sum_{j=1}^{n_{0}}a_{0j}x\sigma^{-m}(a_{mn_{m}})\sigma^{-m}(r)b_{0j}-\sum_{j=1}^{n_{0}}a_{mn_{m}}ra_{0j}xb_{0j}$
$=\phi\big{(}\sum_{j=1}^{n_{0}}a_{0j}\otimes\sigma^{-m}(a_{mn_{m}})\sigma^{-m}(r)b_{0j}-\sum_{j=1}^{n_{0}}a_{mn_{m}}ra_{0j}\otimes b_{0j}\big{)},$
we deduce that $\phi$ is an isomorphism, and
$\sum_{j=1}^{n_{0}}a_{0j}\otimes\sigma^{-m}(a_{mn_{m}})\sigma^{-m}(r)b_{0j}-\sum_{j=1}^{n_{0}}a_{mn_{m}}ra_{0j}\otimes b_{0j}\neq 0,$
so the polynomial
$\sum_{j=1}^{n_{0}}a_{0j}x\sigma^{-m}(a_{mn_{m}})\sigma^{-m}(r)b_{0j}-\sum_{j=1}^{n_{0}}a_{mn_{m}}ra_{0j}xb_{0j}\not\equiv 0.$
Therefore, $g\not\equiv 0$. Furthermore, for all $r\in D$, $g(r)=0$. Thus, the polynomial $g$ is a $\sigma^{m}$-GPI of $D$. We can choose $n_{m}=1$. Then, $g$ is a polynomial not containing $x^{\sigma^{m}}$. Thus, $D$ satisfies a $\sigma^{m}$-GPI not containing $x^{\sigma^{m}}$. According to Corollary 2.3, $D$ satisfies a linear mixed $\sigma^{m}$-GPI of the form
$h=\sum_{i=0}^{k}h_{i}(x^{\sigma^{i}}),$
where $h_{i}(x^{\sigma^{i}})=\sum_{j=1}^{n_{k}}c_{ij}x^{\sigma^{i}}d_{ij}$ and $k\leq(m-1)$. This contradicts the initial assumption. Now, consider the case $n_{0}=0$. Then,
$f=f(x^{\sigma},\cdots,x^{\sigma^{m}})$ $=\sum_{i=0}^{m}g_{i}(x^{\sigma^{i}})$
$=\sum_{j=1}^{n_{1}}a_{1j}x^{\sigma}b_{1j}+\sum_{j=0}^{n_{2}}a_{2j}x^{\sigma^{2}}b_{2j}+\cdots+\sum_{j=0}^{n_{m}}a_{mj}x^{\sigma^{m}}b_{mj}.$
Since $\sigma$ is an anti-automorphism of $D$, for $t\in D$ there exists $r\in D$ such that $\sigma(r)=t$. Therefore, for every $t\in D$,
00 0 $=f(r^{\sigma},\cdots,r^{\sigma^{m}})$
$=\sum_{j=1}^{n_{1}}a_{1j}tb_{1j}+\sum_{j=0}^{n_{2}}a_{2j}t^{\sigma}b_{2j}b_{2j}+\cdots+\sum_{j=0}^{n_{m}}a_{mj}t^{\sigma^{m-1}}b_{mj}.$
Thus, $D$ satisfies a $\sigma^{m}$-GPI of the form
$\sum_{j=1}^{n_{1}}a_{1j}xb_{1j}+\sum_{j=0}^{n_{2}}a_{2j}x^{\sigma}b_{2j}b_{2j}+\cdots+\sum_{j=0}^{n_{m}}a_{mj}x^{\sigma^{m-1}}b_{mj},$
which contradicts the assumption. Therefore, there exists an element $r\in D$ such that $f(r)\neq 0$.
3. Proof of the main result
In this section, we will present the proof of Theorem 1.1. We start with a very classical result. In fact, in the previous section, we occasionally cited these results directly, but as mentioned, the previous section is supplementary. Therefore, for the sake of clarity, we restate these results here.
Lemma 3.1.
Let $D$ be a division ring with center $F$. If $F$ is infinite, then the following assertions are equivalent.
  1. (1)
    $D$ is centrally finite.
  2. (2)
    $D$ satisfies a PI.
  3. (3)
    $D$ satisfies a GPI.
  4. (4)
    $D$ satisfies a GRI.
Proof.
For an original proof the equivalence of (1) and (2), we refer to [9, Theorem 1] and for the implication (4) to (1), see [4]. ∎
The next result is an important intermediate theorem.
Proposition 3.2.
Let $D$ be a division ring with infinite center $F$, and $\sigma$ be an anti-automorphism of $D$. If $D$ satisfies a $\sigma^{m}$-GPI, then $D$ satisfies a GPI. Furthermore, if $D$ satisfies a $\sigma^{m}$-PI, then $D$ satisfies a PI.
Proof.  According to Corollary 2.3, if $D$ satisfies a $\sigma^{m}$-GPI, then we can assume that $D$ satisfies a $\sigma^{m}$-linear polynomial identity
$f(x_{1},x_{2},\cdots,x_{n},x_{1}^{\sigma},x_{2}^{\sigma},\cdots,x_{n}^{\sigma},\cdots,x_{1}^{\sigma^{m}},x_{2}^{\sigma^{m}},\cdots,x_{n}^{\sigma^{m}})\in D\left\langle X_{m}\right\rangle.$
Then, for all $r_{1},r_{2},\cdots,r_{n}\in D$, we have
$f(r_{1},r_{2},\cdots,r_{n},r_{1}^{\sigma},r_{2}^{\sigma},\cdots,r_{n}^{\sigma},\cdots,r_{1}^{\sigma^{m}},r_{2}^{\sigma^{m}},\cdots,r_{n}^{\sigma^{m}})=0.$
Let
$f_{1}(x_{1},x_{1}^{\sigma},\cdots,x_{1}^{\sigma^{m}})=f(x_{1},r_{2},\cdots,r_{n},x_{1}^{\sigma},r_{2}^{\sigma},\cdots,r_{n}^{\sigma},\cdots,x_{1}^{\sigma^{m}},r_{2}^{\sigma^{m}},\cdots,r_{n}^{\sigma^{m}}).$
If $f_{1}(x_{1},x_{1}^{\sigma},\cdots,x_{1}^{\sigma^{m}})\neq 0$, then by using Lemma 2.4, $\dim_{F}D<\infty$. Therefore, according to Lemma 3.1, the division ring $D$ satisfies a GPI. Suppose $f_{1}(x_{1},x_{1}^{\sigma},\cdots,x_{1}^{\sigma^{m}})=0$. For all $r_{1},r_{2},\cdots,r_{n},r_{11},r_{12},\cdots,r_{1m}\in D$, we have
$f(r_{1},r_{2},\cdots,r_{n},r_{11},r_{2}^{\sigma},\cdots,r_{n}^{\sigma},\cdots,r_{1m},r_{2}^{\sigma^{m}},\cdots,r_{n}^{\sigma^{m}})=0.$
Next, let
$f_{2}(x_{2},x_{2}^{\sigma},\cdots,x_{2}^{\sigma^{m}})$
$=f(r_{1},x_{2},r_{3}\cdots,r_{n},r_{11},x_{2}^{\sigma},r_{3}^{\sigma},\cdots,r_{n}^{\sigma},\cdots,r_{1m},x_{2}^{\sigma^{m}},r_{3}^{\sigma^{m}},\cdots,r_{n}^{\sigma^{m}}).$
If $f_{2}(x_{2},x_{2}^{\sigma},\cdots,x_{2}^{\sigma^{m}})\neq 0$, then $D$ satisfies a GPI. Conversely, if
$f_{2}(x_{2},x_{2}^{\sigma},\cdots,x_{2}^{\sigma^{m}})=0$
then for all $r_{1},r_{2},\cdots,r_{n},r_{11},r_{12},\cdots,r_{1m},r_{21},r_{22},\cdots,r_{2m}\in D$, we have
$f(r_{1},r_{2},\cdots,r_{n},r_{11},r_{21},\cdots,r_{n}^{\sigma},\cdots,r_{1m},r_{2m},\cdots,r_{n}^{\sigma^{m}})=0.$
By repeating the same method for $x_{3},x_{4},\cdots,x_{n}$, we can assume that for all $r_{1},r_{2},\cdots,r_{nm}\in D$,
$f(r_{1},r_{2},\cdots,r_{11},r_{21},\cdots,r_{n1},\cdots,r_{1m},r_{2m},\cdots,r_{nm})=0.$
Thus, $D$ satisfies a GPI.
To show the main result of this paper, we borrow the following well known property.
Lemma 3.3.
Let $F$ be a field and $R$ be a ring whose center contains $F$. If $f(t)=a_{n}t^{n}+a_{n-1}t^{n-1}+\dots+a_{1}t+a_{0}$ is a nonzero polynomial over $R$, then $f(t)=0$ has at most $n$ solutions in $F$.
Proof.
This lemma is a corollary of [13, Proposition 2.3.27]. ∎
Now we are ready to prove the main theorem.
The proof of Theorem 1.1.
Let
$f=f(x_{1},\cdots,x_{n},x_{1}^{\sigma},\cdots,x_{n}^{\sigma},\cdots,x_{1}^{\sigma^{m}},\cdots,x_{n}^{\sigma^{m}})$
be a $\sigma^{m}$-GRI of $D$ and $f$ is defined at
$(r_{1},\cdots,r_{n},r_{1}^{\sigma},\cdots,r_{n}^{\sigma},\cdots,r_{1}^{\sigma^{m}},\cdots,r_{n}^{\sigma^{m}}).$
Assume that $t$ is a commuting indeterminate over $D$ and $t^{\sigma}=t$. Then,
$(r_{j}+x_{j}t)^{\sigma^{i}}=r_{j}^{\sigma^{i}}+x_{j}^{\sigma^{i}}t.$
According to [5, Lemma 7], $f$ is also defined at
$(r_{1}+x_{1}t,\cdots,r_{n}+x_{n}t,r_{1}^{\sigma}+x_{1}^{\sigma}t,\cdots,r_{n}^{\sigma}+x_{n}^{\sigma}t,\cdots,r_{1}^{\sigma^{m}}+x_{1}^{\sigma^{m}}t,\cdots,r_{n}^{\sigma^{m}}+x_{n}^{\sigma^{m}}t)$
which is in $D(X_{m})((t))^{n}$. Here $D(X_{m})((t))^{n}$ is the ring of Laurent series over $D(X_{m})$. Furthermore, we can write $f$ as
$f(r_{1}+x_{1}t,\cdots,r_{n}+x_{n}t,r_{1}^{\sigma}+x_{1}^{\sigma}t,\cdots,r_{n}^{\sigma}+x_{n}^{\sigma}t,\cdots,r_{1}^{\sigma^{m}}+x_{1}^{\sigma^{m}}t,\cdots,r_{n}^{\sigma^{m}}+x_{n}^{\sigma^{m}}t)$
$=f_{0}+\sum_{i=1}^{\infty}f_{i}(x_{1},\cdots,x_{n},x_{1}^{\sigma},\cdots,x_{n}^{\sigma},\cdots,x_{1}^{\sigma^{m}},\cdots,x_{n}^{\sigma^{m}})t^{i}$
where $f_{0}=f(r_{1},\cdots,r_{n},r_{1}^{\sigma},\cdots,r_{n}^{\sigma},\cdots,r_{1}^{\sigma^{m}},\cdots,r_{n}^{\sigma^{m}})$ and $f_{i}$ $(i\geq 1)$ are homogeneous polynomials of degree $i$ in $D\langle X_{m}\rangle$. Moreover, there exists $i_{0}$ such that $f_{i_{0}}\neq 0$. We have
$f(r_{1}+d_{1}t,\cdots,r_{n}+d_{n}t,r_{1}^{\sigma}+d_{1}^{\sigma}t,\cdots,r_{n}^{\sigma}+d_{n}^{\sigma}t,\cdots,r_{1}^{\sigma^{m}}+d_{1}^{\sigma^{m}}t,\cdots,r_{n}^{\sigma^{m}}+d_{n}^{\sigma^{m}}t)$
$=\sum_{i=1}^{\infty}f_{i}(d_{1},\cdots,d_{n},d_{1}^{\sigma},\cdots,d_{n}^{\sigma},\cdots,d_{1}^{\sigma^{m}},\cdots,d_{n}^{\sigma^{m}})t^{i}.$
Since $f(r_{1}+d_{1}t,\cdots,r_{n}+d_{n}t,r_{1}^{\sigma}+d_{1}^{\sigma}t,\cdots,r_{n}^{\sigma}+d_{n}^{\sigma}t,\cdots,r_{1}^{\sigma^{m}}+d_{1}^{\sigma^{m}}t,\cdots,r_{n}^{\sigma^{m}}+d_{n}^{\sigma^{m}}t)$ is an element in the quotient ring $D(t)$ of the polynomial ring $D[t]$, it can be represented as
$f(r_{1}+d_{1}t,\cdots,r_{n}+d_{n}t,r_{1}^{\sigma}+d_{1}^{\sigma}t,\cdots,r_{n}^{\sigma}+d_{n}^{\sigma}t,\\ \cdots,r_{1}^{\sigma^{m}}+d_{1}^{\sigma^{m}}t,\cdots,r_{n}^{\sigma^{m}}+d_{n}^{\sigma^{m}}t)=\frac{g(t)}{h(t)}$
where $g(t),h(t)\in D[t]$ and $h(t)\neq 0$. Since the field $F$ is infinite, there exist infinitely many $p\in F$ such that $f$ is defined at $(r_{1}+d_{1}p,\cdots,r_{n}+d_{n}p)$. Furthermore, since $f$ is a $\sigma^{m}$-GRI of $D$, we have
$f(r_{1}+d_{1}p,\cdots,r_{n}+d_{n}p,r_{1}^{\sigma}+d_{1}^{\sigma}p,\cdots,r_{n}^{\sigma}+d_{n}^{\sigma}p,\cdots,r_{1}^{\sigma^{m}}+d_{1}^{\sigma^{m}}p,\cdots,r_{n}^{\sigma^{m}}+d_{n}^{\sigma^{m}}p)$
$=\frac{g(p)}{h(p)}=0,$
implying $g(p)=0$. Thus, the polynomial $g(t)$ has infinitely many roots in $F$, so $g(t)\equiv 0$ by Lemma 3.3. In this case,
$f(r_{1}+d_{1}t,\cdots,r_{n}+d_{n}t,r_{1}^{\sigma}+d_{1}^{\sigma}t,\cdots,r_{n}^{\sigma}+d_{n}^{\sigma}t,\cdots,r_{1}^{\sigma^{m}}+d_{1}^{\sigma^{m}}t,\cdots,r_{n}^{\sigma^{m}}+d_{n}^{\sigma^{m}}t)=0,$
hence
$=\sum_{i=1}^{\infty}f_{i}(d_{1},\cdots,d_{n},d_{1}^{\sigma},\cdots,d_{n}^{\sigma},\cdots,d_{1}^{\sigma^{m}},\cdots,d_{n}^{\sigma^{m}})t^{i}=0.$
Therefore, $f_{i}(d_{1},d_{2},\cdots,d_{n},d_{1}^{\sigma},d_{2}^{\sigma},\cdots,d_{n}^{\sigma},\cdots,d_{1}^{\sigma^{m}},d_{2}^{\sigma^{m}},\cdots,d_{n}^{\sigma^{m}})=0$ for all $i$. Consequently, for all $d_{1},d_{2},\cdots,d_{n}\in D$, we have
$f_{i_{0}}(d_{1},d_{2},\cdots,d_{n},d_{1}^{\sigma},d_{2}^{\sigma},\cdots,d_{n}^{\sigma},\cdots,d_{1}^{\sigma^{m}},d_{2}^{\sigma^{m}},\cdots,d_{n}^{\sigma^{m}})=0.$
Thus,
$f_{i_{0}}(x_{1},x_{2},\cdots,x_{n},x_{1}^{\sigma},x_{2}^{\sigma},\cdots,x_{n}^{\sigma},\cdots,x_{1}^{\sigma^{m}},x_{2}^{\sigma^{m}},\cdots,x_{n}^{\sigma^{m}})$
is a $\sigma^{m}$-$GPI$ of $D$. According to Proposition 3.2, the ring $D$ satisfies a GPI. Therefore, $D$ is centrally finite by Lemma 3.1. ∎